OpenStudy (brittanie.hair):

Someone help me understand how to determine the end behavior of the graph of a polynomial function. Explain using this polynomial function y= 5x^3-2x^2+1

OpenStudy (brittanie.hair):

@satellite73

OpenStudy (anonymous):

the degree is 3 right?

OpenStudy (anonymous):

and three is odd

OpenStudy (anonymous):

the leading coefficient is positive, ( it is 5)

OpenStudy (anonymous):

those two facts taken together means \[\lim_{x\to -\infty}f(x)=-\infty\\ \lim_{x\to \infty}f(x)=\infty\]

OpenStudy (anonymous):



OpenStudy (brittanie.hair):

Oh okay so it would just be an odd degree

OpenStudy (anonymous):

it is an odd degree, it is 3 the same would be true if the degree was five

OpenStudy (brittanie.hair):

Alright I think I get it. Can we try another from my homework?

OpenStudy (anonymous):

oops computer crash yes , if you are stillthere